OpenStudy (exoticinspirit):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)? I am doing a review I know the equation is (h-k)^2+(y-k)^2=r^2 However, I forgot how to plug in the coordinates. Can anyone help me out please? I appreciate it.

OpenStudy (exoticinspirit):

@dan815

OpenStudy (exoticinspirit):

@amistre64

OpenStudy (amistre64):

your 'knowing' is a little off

OpenStudy (amistre64):

for some center (h,k) and radius r .... (x-h)^2+(y-k)^2=r^2

OpenStudy (amistre64):

the radius of a circle is the distance from the center to a point in the edge .... how do we find the distance between 2 points?

OpenStudy (amistre64):

well, we just use the circle equation really a point (a,b) on a circle satisfies (a-h)^2 + (b-k)^2 = r^2

OpenStudy (exoticinspirit):

Thanks I already found the answer though